OpenStudy (anonymous):

Janie's couch is 7 1/2 feet long.The scale of the dollhouse is 1 inch: 2 feet. What is the length of the miniature couch for the dollhouse?

OpenStudy (studygurl14):

You can set up a proportion. 7 1/2 --> 7.5 (just to make it easier...we'll convert back to fraction later) Scale: 1 inch to 2 feet --> \(\Large\frac{1in}{2ft}\) You want to find out how many inches correlates to the actual length of the couch. \(\Large\rightarrow\frac{x\ in}{7.5ft}=\frac{1in}{2ft}\)

OpenStudy (studygurl14):

You want to solve for x in the equation above. Can you do the rest yourself?

OpenStudy (studygurl14):

It's not 2 1/2 Cross multiply to solve: \(\Large\frac{x\ in}{7.5ft}=\frac{1in}{2ft}\rightarrow 2x=7.5(1)\rightarrow 2x=7.5\) Solve for x

OpenStudy (studygurl14):

3.75 or 3 3/4 inches is correct

OpenStudy (studygurl14):

i think you mean 7.5 divided by 2. but yes, 7.5/2 = 3.75

OpenStudy (studygurl14):

I don't know what you mean...3.75 = 3 3/4

OpenStudy (anonymous):

good work

OpenStudy (anonymous):

this is great @codilee98